OpenStudy (anonymous):

For the function g(x)=x^5-16x^3 solve the following. g(x)>0

OpenStudy (anonymous):

Can you factor anything out from each term?

OpenStudy (bahrom7893):

trickets u never responded to mkuehn. Look at what is similar in each terms.

OpenStudy (anonymous):

x can be factored out?

OpenStudy (bahrom7893):

not just x, x to some power.

OpenStudy (anonymous):

If you can factor out x, try factoring out x^2. Can you factor out even more than x^2 ?

OpenStudy (anonymous):

factor out x^3? which gives me x^2-16?

OpenStudy (anonymous):

Well it gives you this: (x^3)(x^2 - 16) > 0

OpenStudy (anonymous):

So now we basically want to find the zeros of this function. Can you factor anything else in order to figure out the zeros of this function?

OpenStudy (bahrom7893):

true so now factor the other side (hint a^2-b^2 = (a+b)(a-b) )

OpenStudy (anonymous):

sorry guys, im just not getting this... so my equation now is (x^3)(x^2-16)>0 ?

OpenStudy (anonymous):

Yes can you factor the x^2 - 16 ?

OpenStudy (anonymous):

If you can imagine the graph of this function, it is a cubic, so it goes up and down and up and down along the x-axis, we are trying to figure out all the places where it crosses the x-axis and then we can determine which intervals it is positive.

OpenStudy (bahrom7893):

rewrite the x^2-16 as x^2-4^2 Then use: a^2-b^2=(a-b)(a+b)

OpenStudy (bahrom7893):

where a=x and b=4

OpenStudy (bahrom7893):

okay i think mkuehn is doing a good job, i will let him finish

OpenStudy (anonymous):

so (x-4) and (x+4)?

OpenStudy (anonymous):

Yes! Exactly!

OpenStudy (anonymous):

So now to find the zeros we have (x^3)(x-4)(x+4) Set each of those factors equal to zero and solve for x.

OpenStudy (anonymous):

x^3 = 0 x = ? x + 4 = 0 x = ? x - 4 = 0 x = ?

OpenStudy (anonymous):

x=0? on all of them?

OpenStudy (anonymous):

wait no 0, -4, 4?

OpenStudy (anonymous):

It works for the first one. If x = 0 for the second one then 0 + 4 = 0 ? Does that make sense? What value of x + 4 equals 0?

OpenStudy (anonymous):

You are answering the question what number plus four equals zero ?

OpenStudy (anonymous):

0, -4, and 4?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So now we know this is where our graph crosses the x-axis, so we basically have 4 intervals we need to look at to see if the function is positive or negative.

OpenStudy (anonymous):

From negative infinity to -4 From -4 to 0 From 0 to 4 From 4 to infinity

OpenStudy (anonymous):

ok im starting to follow along..

OpenStudy (anonymous):

So what we can do here is pick a value inside of each interval and evaluate the function. If the value we get back is positive then we know g(x) is greater than 0 inside that interval. If we get a value that is negative then we know g(x) is less than 0 inside that interval.

OpenStudy (anonymous):

so (-oo,-4)U(4,oo)?

OpenStudy (anonymous):

I'll show you the first example. For the interval from negative infinity to -4 I will select -5. g(-5) = (-5)^5 - 16(-5)^3 = -3125 - 16(-125) = -3125 + 2000 = -1125 So from negative infinity to -4 the function is less than 0. We don't want that, so that is not part of the solution.

OpenStudy (anonymous):

Now check a value between -4 and 0 to check to see if it is positive or negative. You are on the right track.

OpenStudy (bahrom7893):

Great job mkuehn10!

OpenStudy (bahrom7893):

u too tricketts8737

OpenStudy (anonymous):

positive?

OpenStudy (anonymous):

Check out this picture of the function for reference. YES IT IS POSITIVE, so the interval from -4 to 0 is part of the solution.

OpenStudy (anonymous):

Hmm that graph doesn't look right.

OpenStudy (zarkon):

you have -x^5

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

There, that should make it a lot clearer.

OpenStudy (anonymous):

(-oo,-4)U(0,4)?

OpenStudy (anonymous):

So as we figured out it is negative from negative infinity to -4 and positive from -4 to 0. We just need to figure out from 0 to 4 and 4 to infinity. Can you tell by looking at the graph?

OpenStudy (anonymous):

That interval is where it is less than zero.

OpenStudy (anonymous):

We want greater than zero. You are right on the cusp of getting it.

OpenStudy (anonymous):

(-4, 0) U (0,oo)

OpenStudy (anonymous):

You can see this on the graph I attached as well as proving it by checking values inside the intervals.

OpenStudy (anonymous):

ahhh... ok a visual graph would help me alot, i just have a hard time understanding this!

OpenStudy (anonymous):

Why are the zeros not included in the intervals?

OpenStudy (anonymous):

Did you see the file I attached? I will attach it to this message too.

OpenStudy (anonymous):

yes! i see it

OpenStudy (anonymous):

Pretty easy to see where it is negative and positive there. And we showed it with the math.

OpenStudy (anonymous):

so that makes it (-oo,-4)U(4,oo)?

OpenStudy (anonymous):

If you go from left to right on the graph you see it is negative from -infinity to -4 then it is positive from -4 to 0 negative again from 0 to 4 and positive from 4 to infinity

OpenStudy (anonymous):

Since we want where the function is greater than zero, it corresponds with the intervals where the function is positive. From -4 to 0 and from 4 to infinty (-4, 0) U (4, oo)

OpenStudy (anonymous):

ohh ok that makes since.. now that i put the visual concept behind it!! thank you so much!

OpenStudy (anonymous):

You're welcome!